IPP Role in CDF

IPP effort has been focused on t and b physics

• Involved in top quark search and kinematic analysis

• Now using top signal to search for new physics – pp̄ → W ′ → tb̄X

– complements another IPP analysis effort (W ′ → e±νe)

• Have five separate b analysis projects 1. Search for B → K∗γ

2. BR of Bd, Bu and Bs into J/ψK, J/ψK∗ and J/ψϕ 3. First high-statistics study of B → ψ(2S)K and ψ(2S)K∗ 4. b quark fragmentation into hadrons

5. Study of pp̄ → χcX

For Run 2, will continue to focus on b and t physics

• Search for rare decays

• b-flavour tagging for CP studies

• High-statistics top studies

IPP Responsibilities in CDF

IPP has visible role in collaboration

• Managed Offline Software and Data Processing group • Contributed to online software and firmware

• 2 FTE’s involved in software development/maintenance • R&D in Level 1 tracking trigger and fast RICH

1. Have developed prototype ASIC’s for L1 trigger

2. Demonstrating fast RICH proof of principle

With new member (W. Trischuk) joining group

• Are evaluating options for Run 2 responsibilities

• Most likely involvement is in tracking upgrade – builds on DELPHI experience

– possibility of replacing inner SVX layers with diamond

sensors

• Will have made decision in next 6 months

Resource Commitments

Personnel involved in CDF effort is increasing

95-96  96-97  97-98  98-99  99-00
Ragan  1.0  1.0  1.0  1.0  0.7
Sinervo  0.9  0.9  0.9  0.8  0.7
Trischuk  0.9  0.9  0.8  0.7
Total Faculty  1.9  2.8  2.8  2.6  2.1
RA  2.0  2.0  2.0  2.0  2.0
Students  6  7  8  8  8
Engineering  0.5  0.5  0.5  0.3  0.1
Total Effort  10.4  12.3  13.3  12.9  12.2

ATLAS is future project for two of us (Trischuk and Sinervo)

• focus for next two years will remain CDF

• forsee increasing role in ATLAS after 1997

Budget Projections

Budget based on existing operating support level (193K)

• require increment to support Trischuk’s activity

• require equipment & computing support – IPP is required to make significant Run 2 investment – One example: 30-50% of CVD Diamond upgrade

∗ commitment would be ∼ 300K over 97-98 and 98-99

95-96  96-97  97-98  98-99  99-00
Operating  193  300  300  250  250
Equipment  0  0  150  150  0
Computing  0  100  0  0  100
Total Budget  193  400  450  400  350

Other Funding Scenarios

Have considered the scenarios of constant funding, ±10%

• All of these are inconsistent with a Run 2 effort – Have increased FTE effort by ∼ 50%

– Require equipment contribution to Run 2 detector – Computer facilities inadequate for Run 1 analysis

Only reasonable options are

1. Adequately fund a Run 2 effort

2. Turn IPP effort on CDF off over next 3 years

We believe that option 1) is most productive

• Some of the most exciting physics in year 2000

• Additional investment is modest and “strategic”

• Excellent background for future collider programme